### v1

@@ 2 Preliminaries
Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\bm{W}})$ be an undirected weighted graph on $n=|{\mathcal{V}}|$ vertices with a symmetric, elementwise nonnegative adjacency matrix ${\bm{W}}\in\mathbb{R}^{n\times n}$; assume ${\bm{W}}_{ii}=0$. Define the degree of $i\in{\mathcal{V}}$ by $d_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{j\in{\mathcal{V}}}{\bm{W}}_{ij}$ and the degree matrix ${\bm{D}}\stackrel{{\scriptstyle\text{def}}}{{=}}\mathop{\mathrm{diag}}(d_{1},\ldots,d_{n})$.
For ${\mathbb{A}}\subseteq{\mathcal{V}}$, let $\overline{{\mathbb{A}}}\stackrel{{\scriptstyle\text{def}}}{{=}}{\mathcal{V}}\setminus{\mathbb{A}}$ and identify $\mathbb{A}$} by its indicator vector ${\bm{1}}_{{\mathbb{A}}}\in\{0,1\}^{n}$. The cut associated with $\mathbb{A}$} is:
$\operatorname*{cut}({\mathbb{A}},\overline{{\mathbb{A}}})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{(v_{i},v_{j})\in{\mathbb{A}}\times\overline{{\mathbb{A}}}}{\bm{W}}_{ij}\;=\;{\bm{1}}_{{\mathbb{A}}}^{\top}{\bm{W}}\,{\bm{1}}_{\overline{{\mathbb{A}}}},$
@@
$\operatorname*{VolCut}({\mathbb{A}})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{\operatorname*{cut}({\mathbb{A}},\overline{{\mathbb{A}}})}{\operatorname*{vol}({\mathbb{A}})},$ (1)
where the volume is $\operatorname*{vol}({\mathbb{A}})\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{v_{i}\in{\mathbb{A}}}s(v_{i})$ for a given vertex weight function $s:{\mathcal{V}}\to\mathbb{R}_{>0}$. For example, the ratio cut uses $s(v_{i})\equiv 1$ so $\operatorname*{vol}\left({\mathbb{A}}\right)\equiv|{\mathbb{A}}|$, whereas the normalized cut uses $s(v_{i})=d_{i}$.
Fix $s$ and write $s_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}s(v_{i})$. The goal is to find a $k$-way partition ${\mathcal{C}}_{k}=\{{\mathbb{C}}_{\ell}\}_{\ell=1}^{k}$ of $\mathcal{V}$} that minimizes the volume-normalized graph cut:
$\text{GraphCut}({\mathcal{C}}_{k})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{1}{2}\sum_{\ell=1}^{k}\operatorname*{VolCut}({\mathbb{C}}_{\ell}).$ (2)
@@ 2.1 Probabilistic Relaxation
$\text{GraphCut}({\bm{P}})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{1}{2}\sum_{\ell=1}^{k}\mathbb{E}\left[\widehat{\operatorname*{VolCut}}({\mathbf{a}}_{\ell})\right],$ (3)
@@
$\widehat{\operatorname*{VolCut}}({\mathbf{a}}_{\ell})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{{\mathbf{a}}_{\ell}^{\top}{\bm{W}}\,({\bm{1}}-{\mathbf{a}}_{\ell})}{\sum_{i=1}^{n}s_{i}\,{\textnormal{a}}_{\ell,i}}.$ (4)
@@
where $\overline{{\bm{P}}}_{:,\ell}\stackrel{{\scriptstyle\text{def}}}{{=}}\tfrac{1}{n}\sum_{i=1}^{n}{\bm{P}}_{i\ell}$ denotes the expected fraction of vertices assigned to $\mathbb{C}$$_{\ell}$.
@@ 3 Proposed methods
In our setting, $m\stackrel{{\scriptstyle\text{def}}}{{=}}n-2$, ${\bm{\alpha}}=(p_{3},\ldots,p_{n})$, and the weights are $\theta_{i}=0$ and $\beta_{i}=s_{i}$, so ${\textnormal{x}}=\sum_{i=3}^{n}s_{i}\,{\textnormal{r}}_{i}$. We denote this special case by $\operatorname{GPB}({\bm{\alpha}},{\bm{\beta}})$, and compute its probability generating function (PGF) $G_{\textnormal{x}}$:
$G_{\textnormal{x}}(t)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}\left[t^{{\textnormal{x}}}\right]=\prod_{i=1}^{m}\big(1-\alpha_{i}+\alpha_{i}\,t^{\,\beta_{i}}\big),\;t\in[0,1].$ (6)
@@
$\mathbb{E}\left[\frac{1}{q+{\textnormal{x}}}\right]=\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\int_{0}^{1}t^{\,q-1}\,\prod_{i=1}^{m}\big(1-\alpha_{i}+\alpha_{i}\,t^{\,\beta_{i}}\big)\,dt.$ (7)
@@
where $(x)_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}x(x+1)\cdots(x+k-1)$ and $(x)_{0}\stackrel{{\scriptstyle\text{def}}}{{=}}1$ is the rising factorial.
@@
Let $m\in\mathbb{N}$, $b>0$, and $c>b$. On $[0,1]$, the function $f(z)\stackrel{{\scriptstyle\text{def}}}{{=}}{}_{2}F_{1}(-m,b;c;z)$ is a degree-$m$ polynomial that is decreasing, convex, and $L$-Lipschitz with $L=\tfrac{mb}{c}$.
@@
where $\bar{\alpha}\stackrel{{\scriptstyle\text{def}}}{{=}}{\frac{1}{m}}\sum_{i=1}^{m}\alpha_{i}$. We denote the right-hand side by ${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)$.
@@ 3.1 The AM-GM gap
$\underline{\Delta}(q,{\bm{\alpha}})$ $\stackrel{{\scriptstyle\text{def}}}{{=}}\int_{0}^{1}h(t)\Bigl(1-e^{-\gamma(t)\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
$\overline{\Delta}(q,{\bm{\alpha}})$ $\stackrel{{\scriptstyle\text{def}}}{{=}}\int_{0}^{1}h(t)\Bigl(1-e^{-\theta(t)\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
with $\gamma(t)\stackrel{{\scriptstyle\text{def}}}{{=}}\tfrac{m}{2}(1-t^{\beta})^{2}$ and $\theta(t)\stackrel{{\scriptstyle\text{def}}}{{=}}\gamma(t)/t^{2\beta}$. Then:
@@
By Equation 7, coordinates with $\alpha_{i}=0$ contribute the factor $(1-\alpha_{i}+\alpha_{i}t^{\beta})\equiv 1$ and thus make no contribution to ${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})$. The AM–GM gap in Proposition 3.6 over-penalizes configurations with many inactive entries: zeros still inflate $\mathrm{Var}({\bm{\alpha}})$ even though they do not affect the product inside the integral. We therefore replace the plain variance by a zero-aware weighted dispersion that vanishes at $\alpha_{i}=0$. Let $\omega_{0}(x)\stackrel{{\scriptstyle\text{def}}}{{=}}x$ (more generally, $\omega_{0}(x)=x^{a},a\in[1,2]$), and define:
$\Omega\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\sum_{i=1}^{m}\omega_{0}(\alpha_{i}),\qquad\bar{\alpha}^{\omega_{0}}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{1}{\Omega}\sum_{i=1}^{m}\omega_{0}(\alpha_{i})\,\alpha_{i},$ (13)
$\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{1}{\Omega}\sum_{i=1}^{m}\omega_{0}(\alpha_{i})\bigl(\alpha_{i}-\bar{\alpha}^{\omega_{0}}\bigr)^{2},$ (14)
@@
${\mathcal{A}}(q;\bar{\alpha},m)\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\;{\mathcal{H}}_{\beta}\!\bigl(q+r\beta;\bar{\alpha},m\bigr).$ (15)
@@
${\mathbb{A}}(q,{\bm{\alpha}},m)\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{m}{2}\;\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\widetilde{{\mathcal{A}}}\bigl(q;\bar{\alpha},m\bigr).$ (16)
@@ 3.2 Concentration for minibatch estimators
Fix $m\in\mathbb{N}$, $q>0$, $\beta>0$. Let $c=\tfrac{q}{\beta}$ and $\{\alpha_{i}\}_{i=1}^{m}\subset[0,1]$ be the (fixed) population, with mean $\bar{\alpha}\!=\!m^{-1}\sum_{i}\alpha_{i}$ and population variance $\sigma^{2}\!\stackrel{{\scriptstyle\text{def}}}{{=}}\!m^{-1}\sum_{i}(\alpha_{i}-\bar{\alpha})^{2}$. Form a minibatch $S=(I_{1},\dots,I_{n})$ of size $n$ by sampling indices with replacement, and define the plug-in estimator:
$\hat{\alpha}_{S}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\tfrac{1}{n}\sum_{r=1}^{n}\alpha_{I_{r}},\qquad\tilde{H}(S)\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;{\mathcal{H}}_{\beta}\!\bigl(q;\hat{\alpha}_{S},m\bigr).$ (17)
@@
$\frac{d^{2}}{dz^{2}}\,{\mathcal{H}}_{\beta}(q;z,m)\;\leq\;\underbrace{\frac{1}{q}\cdot\frac{2m(m-1)}{c(c+1)}}_{\!\!\!\stackrel{{\scriptstyle\text{def}}}{{=}}K},$ (19)
@@ 3.3 Heterogeneous degrees
When $(\beta_{i})_{i}$ vary, directly using a single $\beta$ loses heterogeneity. We partition indices into $d$ disjoint bins $S_{1},\dots,S_{d}$ based on their $\beta_{i}$ values. Let $m_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}|S_{j}|$, $\bar{\alpha}_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}m_{j}^{-1}\sum_{i\in S_{j}}\alpha_{i}$, and define the bin interval $B_{j}=[b_{j-1},b_{j}]$ with representatives $\beta_{j}^{\star}\in B_{j}$ specified below.
@@ 3.4 Optimization objective
Fix $\ell\in\{1,\dots,k\}$ and partition indices into $d$ bins $S_{\ell 1},\ldots,S_{\ell d}$ by their exponents $\beta_{u}\equiv s_{u}$ (e.g., degree-based); let $m_{\ell j}\stackrel{{\scriptstyle\text{def}}}{{=}}|S_{\ell j}|$, $m_{\ell}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{j}m_{\ell j}$, and
$\bar{p}_{\ell j}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{1}{m_{\ell j}}\sum_{u\in S_{\ell j}}{\bm{P}}_{u\ell},\qquad w_{\ell j}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{m_{\ell j}}{m_{\ell}}.$
@@
$\Phi_{\ell}(q)\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\prod_{j=1}^{d}\Bigl[\;{\mathcal{H}}_{\beta_{\ell j}^{\star}}\!\bigl(q;\bar{p}_{\ell j},\,m_{\ell}\bigr)\;\Bigr]^{w_{\ell j}}.$
@@
$M_{i\ell}({\bm{P}})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\sum_{j=1}^{n}{\bm{W}}_{ij}\,{\bm{P}}_{i\ell}\,\bigl(1-{\bm{P}}_{j\ell}\bigr),$
@@
$U_{\ell}({\bm{P}})\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=1}^{n}M_{i\ell}({\bm{P}})\,\Phi_{\ell}\!\bigl(s_{i}\bigr),$ (26)
and $U({\bm{P}})\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{\ell=1}^{k}U_{\ell}({\bm{P}})$. By construction (linearity of expectation and Theorem 3.10), $I_{\mathrm{true}}\!\leq U$.
@@
$\Gamma_{\ell}({\bm{P}})\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=1}^{n}M_{i\ell}({\bm{P}})\,\Biggl[\;\sum_{j=1}^{d}w_{\ell j}\,{\mathbb{A}}(\beta_{\ell j}^{\star},{\bm{p}}_{\ell j},m_{\ell})\Biggr],$ (27)
and $\mathbb{A}$} is the zero–aware coefficient fromProposition 3.8. Summing over clusters, $\Gamma({\bm{P}})\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{\ell=1}^{k}\Gamma_{\ell}({\bm{P}})$ satisfies
@@
$\boxed{\min_{\;{\bm{P}}={\bm{P}}({\bm{z}})}{\mathbb{J}}_{\rho}({\bm{P}})\stackrel{{\scriptstyle\text{def}}}{{=}}U({\bm{P}})\;+\;\rho\,\Gamma({\bm{P}}),\quad\rho\geq 0,\quad}$ (28)

### v2

@@ Abstract
Probabilistic Graph Cuts
Ayoub Ghriss
Department of Computer Science
University of Colorado, Boulder
ayoub.ghriss@colorado.edu
@@ 2 Preliminaries
Let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\bm{W}})$ be an undirected weighted graph on $n=|{\mathcal{V}}|$ vertices with a symmetric, elementwise nonnegative adjacency matrix ${\bm{W}}\in\mathbb{R}^{n\times n}$; assume ${\bm{W}}_{ii}=0$. Define the degree of $i\in{\mathcal{V}}$ by $d_{i}\vcentcolon=\sum_{j\in{\mathcal{V}}}{\bm{W}}_{ij}$ and the degree matrix ${\bm{D}}\vcentcolon=\mathop{\mathrm{diag}}(d_{1},\ldots,d_{n})$.
For ${\mathbb{A}}\subseteq{\mathcal{V}}$, let $\overline{{\mathbb{A}}}\vcentcolon={\mathcal{V}}\setminus{\mathbb{A}}$ and identify $\mathbb{A}$} by its indicator vector ${\bm{1}}_{{\mathbb{A}}}\in\{0,1\}^{n}$. The cut associated with $\mathbb{A}$} is:
$\operatorname*{cut}({\mathbb{A}},\overline{{\mathbb{A}}})\;\vcentcolon=\sum_{(v_{i},v_{j})\in{\mathbb{A}}\times\overline{{\mathbb{A}}}}{\bm{W}}_{ij}\;=\;{\bm{1}}_{{\mathbb{A}}}^{\top}{\bm{W}}\,{\bm{1}}_{\overline{{\mathbb{A}}}},$
@@
$\operatorname*{VolCut}({\mathbb{A}})\;\vcentcolon=\;\frac{\operatorname*{cut}({\mathbb{A}},\overline{{\mathbb{A}}})}{\operatorname*{vol}({\mathbb{A}})},$ (1)
where the volume is $\operatorname*{vol}({\mathbb{A}})\vcentcolon=\sum_{v_{i}\in{\mathbb{A}}}s(v_{i})$ for a given vertex weight function $s:{\mathcal{V}}\to\mathbb{R}_{>0}$. For example, the ratio cut uses $s(v_{i})\equiv 1$ so $\operatorname*{vol}\left({\mathbb{A}}\right)\equiv|{\mathbb{A}}|$, whereas the normalized cut uses $s(v_{i})=d_{i}$.
Fix $s$ and write $s_{i}\vcentcolon=s(v_{i})$. The goal is to find a $k$-way partition ${\mathcal{C}}_{k}=\{{\mathbb{C}}_{\ell}\}_{\ell=1}^{k}$ of $\mathcal{V}$} that minimizes the volume-normalized graph cut:
$\text{GraphCut}({\mathcal{C}}_{k})\;\vcentcolon=\;\frac{1}{2}\sum_{\ell=1}^{k}\operatorname*{VolCut}({\mathbb{C}}_{\ell}).$ (2)
@@ 2.1 Probabilistic Relaxation
$\text{GraphCut}({\bm{P}})\;\vcentcolon=\;\frac{1}{2}\sum_{\ell=1}^{k}\mathbb{E}\left[\widehat{\operatorname*{VolCut}}({\mathbf{a}}_{\ell})\right],$ (3)
@@
$\widehat{\operatorname*{VolCut}}({\mathbf{a}}_{\ell})\;\vcentcolon=\;\frac{{\mathbf{a}}_{\ell}^{\top}{\bm{W}}\,({\bm{1}}-{\mathbf{a}}_{\ell})}{\sum_{i=1}^{n}s_{i}\,{\textnormal{a}}_{\ell,i}}.$ (4)
@@
where $\overline{{\bm{P}}}_{:,\ell}\vcentcolon=\tfrac{1}{n}\sum_{i=1}^{n}{\bm{P}}_{i\ell}$ denotes the expected fraction of vertices assigned to $\mathbb{C}$$_{\ell}$.
@@ 3 Proposed methods
In our setting, $m\vcentcolon=n-2$, ${\bm{\alpha}}=(p_{3},\ldots,p_{n})$, and the weights are $\theta_{i}=0$ and $\beta_{i}=s_{i}$, so ${\textnormal{x}}=\sum_{i=3}^{n}s_{i}\,{\textnormal{r}}_{i}$. We denote this special case by $\operatorname{GPB}({\bm{\alpha}},{\bm{\beta}})$, and compute its probability generating function (PGF) $G_{\textnormal{x}}$:
$G_{\textnormal{x}}(t)\vcentcolon=\mathbb{E}\left[t^{{\textnormal{x}}}\right]=\prod_{i=1}^{m}\big(1-\alpha_{i}+\alpha_{i}\,t^{\,\beta_{i}}\big),\;t\in[0,1].$ (6)
@@
$\mathbb{E}\left[\frac{1}{q+{\textnormal{x}}}\right]=\;\vcentcolon=\;\int_{0}^{1}t^{\,q-1}\,\prod_{i=1}^{m}\big(1-\alpha_{i}+\alpha_{i}\,t^{\,\beta_{i}}\big)\,dt.$ (7)
@@
where $(x)_{k}\vcentcolon=x(x+1)\cdots(x+k-1)$ and $(x)_{0}\vcentcolon=1$ is the rising factorial.
@@
Let $m\in\mathbb{N}$, $b>0$, and $c>b$. On $[0,1]$, the function $f(z)\vcentcolon={}_{2}F_{1}(-m,b;c;z)$ is a degree-$m$ polynomial that is decreasing, convex, and $L$-Lipschitz with $L=\tfrac{mb}{c}$.
@@
where $\bar{\alpha}\vcentcolon={\frac{1}{m}}\sum_{i=1}^{m}\alpha_{i}$. We denote the right-hand side by ${\mathcal{H}}_{\beta}(q;\bar{\alpha},m)$.
@@ 3.1 The AM-GM gap
$\underline{\Delta}(q,{\bm{\alpha}})$ $\vcentcolon=\int_{0}^{1}h(t)\Bigl(1-e^{-\gamma(t)\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
$\overline{\Delta}(q,{\bm{\alpha}})$ $\vcentcolon=\int_{0}^{1}h(t)\Bigl(1-e^{-\theta(t)\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
with $\gamma(t)\vcentcolon=\tfrac{m}{2}(1-t^{\beta})^{2}$ and $\theta(t)\vcentcolon=\gamma(t)/t^{2\beta}$. Then:
@@
By Equation 7, coordinates with $\alpha_{i}=0$ contribute the factor $(1-\alpha_{i}+\alpha_{i}t^{\beta})\equiv 1$ and thus make no contribution to ${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})$. The AM–GM gap in Proposition 3.6 over-penalizes configurations with many inactive entries: zeros still inflate $\mathrm{Var}({\bm{\alpha}})$ even though they do not affect the product inside the integral. We therefore replace the plain variance by a zero-aware weighted dispersion that vanishes at $\alpha_{i}=0$. Let $\omega_{0}(x)\vcentcolon=x$ (more generally, $\omega_{0}(x)=x^{a},a\in[1,2]$), and define:
$\Omega\;\vcentcolon=\;\sum_{i=1}^{m}\omega_{0}(\alpha_{i}),\qquad\bar{\alpha}^{\omega_{0}}\;\vcentcolon=\;\frac{1}{\Omega}\sum_{i=1}^{m}\omega_{0}(\alpha_{i})\,\alpha_{i},$ (13)
$\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\vcentcolon=\;\frac{1}{\Omega}\sum_{i=1}^{m}\omega_{0}(\alpha_{i})\bigl(\alpha_{i}-\bar{\alpha}^{\omega_{0}}\bigr)^{2},$ (14)
@@
${\mathcal{A}}(q;\bar{\alpha},m)\;\vcentcolon=\;\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\;{\mathcal{H}}_{\beta}\!\bigl(q+r\beta;\bar{\alpha},m\bigr).$ (15)
@@
${\mathbb{A}}(q,{\bm{\alpha}},m)\vcentcolon=\frac{m}{2}\;\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\widetilde{{\mathcal{A}}}\bigl(q;\bar{\alpha},m\bigr).$ (16)
@@ 3.2 Concentration for minibatch estimators
Fix $m\in\mathbb{N}$, $q>0$, $\beta>0$. Let $c=\tfrac{q}{\beta}$ and $\{\alpha_{i}\}_{i=1}^{m}\subset[0,1]$ be the (fixed) population, with mean $\bar{\alpha}\!=\!m^{-1}\sum_{i}\alpha_{i}$ and population variance $\sigma^{2}\!\vcentcolon=\!m^{-1}\sum_{i}(\alpha_{i}-\bar{\alpha})^{2}$. Form a minibatch $S=(I_{1},\dots,I_{n})$ of size $n$ by sampling indices with replacement, and define the plug-in estimator:
$\hat{\alpha}_{S}\;\vcentcolon=\;\tfrac{1}{n}\sum_{r=1}^{n}\alpha_{I_{r}},\qquad\tilde{H}(S)\;\vcentcolon=\;{\mathcal{H}}_{\beta}\!\bigl(q;\hat{\alpha}_{S},m\bigr).$ (17)
@@
$\frac{d^{2}}{dz^{2}}\,{\mathcal{H}}_{\beta}(q;z,m)\;\leq\;\underbrace{\frac{1}{q}\cdot\frac{2m(m-1)}{c(c+1)}}_{\!\!\!\vcentcolon=K},$ (19)
@@ 3.3 Heterogeneous degrees
When $(\beta_{i})_{i}$ vary, directly using a single $\beta$ loses heterogeneity. We partition indices into $d$ disjoint bins $S_{1},\dots,S_{d}$ based on their $\beta_{i}$ values. Let $m_{j}\vcentcolon=|S_{j}|$, $\bar{\alpha}_{j}\vcentcolon=m_{j}^{-1}\sum_{i\in S_{j}}\alpha_{i}$, and define the bin interval $B_{j}=[b_{j-1},b_{j}]$ with representatives $\beta_{j}^{\star}\in B_{j}$ specified below.
@@ 3.4 Optimization objective
Fix $\ell\in\{1,\dots,k\}$ and partition indices into $d$ bins $S_{\ell 1},\ldots,S_{\ell d}$ by their exponents $\beta_{u}\equiv s_{u}$ (e.g., degree-based); let $m_{\ell j}\vcentcolon=|S_{\ell j}|$, $m_{\ell}\vcentcolon=\sum_{j}m_{\ell j}$, and
$\bar{p}_{\ell j}\;\vcentcolon=\;\frac{1}{m_{\ell j}}\sum_{u\in S_{\ell j}}{\bm{P}}_{u\ell},\qquad w_{\ell j}\;\vcentcolon=\;\frac{m_{\ell j}}{m_{\ell}}.$
@@
$\Phi_{\ell}(q)\;\vcentcolon=\;\prod_{j=1}^{d}\Bigl[\;{\mathcal{H}}_{\beta_{\ell j}^{\star}}\!\bigl(q;\bar{p}_{\ell j},\,m_{\ell}\bigr)\;\Bigr]^{w_{\ell j}}.$
@@
$M_{i\ell}({\bm{P}})\;\vcentcolon=\;\sum_{j=1}^{n}{\bm{W}}_{ij}\,{\bm{P}}_{i\ell}\,\bigl(1-{\bm{P}}_{j\ell}\bigr),$
@@
$U_{\ell}({\bm{P}})\vcentcolon=\sum_{i=1}^{n}M_{i\ell}({\bm{P}})\,\Phi_{\ell}\!\bigl(s_{i}\bigr),$ (26)
and $U({\bm{P}})\vcentcolon=\sum_{\ell=1}^{k}U_{\ell}({\bm{P}})$. By construction (linearity of expectation and Theorem 3.10), $I_{\mathrm{true}}\!\leq U$.
@@
$\Gamma_{\ell}({\bm{P}})\vcentcolon=\sum_{i=1}^{n}M_{i\ell}({\bm{P}})\,\Biggl[\;\sum_{j=1}^{d}w_{\ell j}\,{\mathbb{A}}(\beta_{\ell j}^{\star},{\bm{p}}_{\ell j},m_{\ell})\Biggr],$ (27)
and $\mathbb{A}$} is the zero–aware coefficient fromProposition 3.8. Summing over clusters, $\Gamma({\bm{P}})\vcentcolon=\sum_{\ell=1}^{k}\Gamma_{\ell}({\bm{P}})$ satisfies
@@
$\boxed{\min_{\;{\bm{P}}={\bm{P}}({\bm{z}})}{\mathbb{J}}_{\rho}({\bm{P}})\vcentcolon=U({\bm{P}})\;+\;\rho\,\Gamma({\bm{P}}),\quad\rho\geq 0,\quad}$ (28)